OpenStudy (anonymous):

Help with solving for x in the equation (2x+3)^2 +x^2 = 40000

OpenStudy (mathmath333):

first u can try simplify this \(\large \begin{align} \color{black}{(2x+3)^2 =\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (anonymous):

4x^2 +12x +9

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{(2x+3)^2 +x^2=40000\hspace{.33em}\\~\\ 4x^2 +12x +9 +x^2=40000\hspace{.33em}\\~\\ 5x^2 +12x +9 =40000\hspace{.33em}\\~\\ x^2 +\dfrac{12x}{5} =\dfrac{39991}{5}\hspace{.33em}\\~\\ x^2 +2\cdot \dfrac{6x}{5} =\dfrac{39991}{5}\hspace{.33em}\\~\\ x^2 +2\cdot \dfrac{6x}{5}+ (\dfrac{6}{5})^2 =\dfrac{39991}{5}+ (\dfrac{6}{5})^2 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ x^2 +2\cdot \dfrac{6x}{5}+ (\dfrac{6}{5})^2 =\dfrac{39991}{5}+ (\dfrac{6}{5})^2 \hspace{.33em}\\~\\ (x+\dfrac{6}{5})^2 =\dfrac{39991}{5}+ \dfrac{36}{25} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Where does dividing by 5 come from?

OpenStudy (mathmath333):

i divided this equation by 5 \(\large \begin{align} \color{black}{ 5x^2 +12x +9 =40000\hspace{.33em}\\~\\ \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

do u know squaring method

OpenStudy (anonymous):

Is it the method used to solve (equation) = 0 equations?

OpenStudy (mathmath333):

yes

OpenStudy (mathmath333):

i wanted to make the red part a perfect squarw \(\large \begin{align} \color{black}{ \color{red}{x^2 +2\cdot \dfrac{6x}{5}+ (\dfrac{6}{5})^2} =\dfrac{39991}{5}+ (\dfrac{6}{5})^2 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

In the first step with dividing by 5, you are getting rid of 5 by dividing both sides of the equation by 5, right?

OpenStudy (anonymous):

Then reducing the 12x to 6x by moving a multiplication of 2 to the left side of the fraction(changing the form but not the value)

OpenStudy (mathmath333):

yes i want to get rid of 5 from the \(x^2\) term

OpenStudy (mathmath333):

yes u r correct!

OpenStudy (anonymous):

What about (6/5)^2 on both sides of the equation?

OpenStudy (mathmath333):

i applied this to red part \(\large \begin{align} \color{black}{ a^2+2ab+b^2=(a+b)^2 \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black}{ x^2 +2\cdot \dfrac{6x}{5}+ (\dfrac{6}{5})^2 =\dfrac{39991}{5}+ (\dfrac{6}{5})^2 \hspace{.33em}\\~\\ (x+\dfrac{6}{5})^2 =\dfrac{199991}{25} \hspace{.33em}\\~\\ (x+\dfrac{6}{5}) =\pm\sqrt{\dfrac{199991}{25}} \hspace{.33em}\\~\\ (x+\dfrac{6}{5}) =\pm\dfrac{1}{5}\sqrt{199991} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (anonymous):

So the format of this equation would take b but not a(removing the fraction but not x) and add it on both sides? Also, how is the left side reduced to \[(x+6/5)^2\]

OpenStudy (mathmath333):

\(\large \begin{align} \color{red}{ x^2 +2\cdot \dfrac{6x}{5}+ (\dfrac{6}{5})^2 \\ =x^2+ \dfrac{6x}{5}+\dfrac{6x}{5}+ (\dfrac{6}{5})^2 \\ =x(x+ \dfrac{6}{5})+\dfrac{6}{5}(x+ (\dfrac{6}{5})) \\ =(x+ \dfrac{6}{5})(x+ (\dfrac{6}{5})) \\ =(x+ \dfrac{6}{5})^2\\ }\end{align}\) see if that makes sense

OpenStudy (anonymous):

So the multiplication of 2 makes 6x/5 become 2 fractions. The x is distributed in both fractions to change the form of everything but what happens to the 6/5 in the middle and the x on the outside?

OpenStudy (mathmath333):

\(\large \begin{align} \color{red}{ x^2 +2\cdot \dfrac{6x}{5}+ (\dfrac{6}{5})^2 \\ =x^2+ \dfrac{6x}{5}+\dfrac{6x}{5}+ (\dfrac{6}{5})^2 \\ =\color{blue}{x}(x+ \dfrac{6}{5})+\color{blue}{\dfrac{6}{5}}(x+ (\dfrac{6}{5})) \\ =(x+ \dfrac{6}{5})(x+ (\dfrac{6}{5})) \\ =(x+ \dfrac{6}{5})^2\\ }\end{align}\) i just took out x and \(6x/5\) common

OpenStudy (anonymous):

Ok i understand now

OpenStudy (mathmath333):

i mean x and 6/5 not 6x/5

OpenStudy (mathmath333):

u know how to proceed from here \((x+\dfrac{6}{5}) =\pm\dfrac{1}{5}\sqrt{199991} \hspace{.33em}\\~\\\)

OpenStudy (mathmath333):

\(\large \begin{align} \color{black} { (x+\dfrac{6}{5}) =\pm\dfrac{1}{5}\sqrt{199991} \hspace{.33em}\\~\\ (x+\dfrac{6}{5}) =+\dfrac{1}{5}\sqrt{199991} \hspace{.33em}\\~\\ \hspace{.33em}\\~\\ or \hspace{.33em}\\~\\ (x+\dfrac{6}{5}) =-\dfrac{1}{5}\sqrt{199991}\hspace{.33em}\\~\\ so \hspace{.33em}\\~\\ x =-\dfrac{1}{5}\sqrt{199991}-\dfrac{6}{5} \hspace{.33em}\\~\\ x =+\dfrac{1}{5}\sqrt{199991}-\dfrac{6}{5} \hspace{.33em}\\~\\ }\end{align}\)

OpenStudy (mathmath333):

check with wolfram http://www.wolframalpha.com/input/?i=solve+%282x%2B3%29%5E2+%2Bx%5E2+%3D+40000

OpenStudy (anonymous):

Ok, i understand it now. Thanks for the help.